OpenStudy (xxdragoonxx):

What is the circumference of a hubcup with a diameter of 10 inches?

OpenStudy (anonymous):

First you need to know that formula of the circumference of a hubcup (also knowns as a circle figure)

OpenStudy (anonymous):

C = 2*pi*r C = 2*3.14*10/2 (Diameter = 10/2 which is 5) C = 2*3.14*5

OpenStudy (anonymous):

C= 31.4

OpenStudy (xxdragoonxx):

Thank you for the explanation.

OpenStudy (anonymous):

Welcome